If c1 and c2 are constant vectors and @ is a constant scalar, show that h=e^(-@x)(c1sin@x+c2cos@y)


1
Expert's answer
2019-09-11T10:41:31-0400

Let's write the condition clearly since the existing condition is incomplete and a little bit confusing:


If "\textbf{c}_1"and "\textbf{c}_2" are constant vectors and "\lambda" is a constant scalar, show that

"\textbf{h}=e^{-\lambda x}[\textbf{c}_1\text{sin}(\lambda x)+\textbf{c}_2 \text{cos}(\lambda y)]."

satisfies the partial equation


"\frac{\partial^2\textbf{h}}{\partial x^2}+\frac{\partial^2\textbf{h}}{\partial y^2}=0."

Solution


Simply take the second partial derivatives of the given vector:


"\Big(\frac{\partial^2\textbf{}}{\partial x^2}+\frac{\partial^2\textbf{}}{\partial y^2}\Big)\textbf{h}=\\
\space\\
=\Big[\frac{\partial^2\textbf{}}{\partial x^2}+\frac{\partial^2\textbf{}}{\partial y^2}\Big]e^{-\lambda x}[\textbf{c}_1\text{sin}(\lambda x)+\textbf{c}_2 \text{cos}(\lambda y)]=\\
\space\\
=\frac{\partial^2\textbf{}}{\partial x^2}e^{-\lambda x}\textbf{c}_1\text{sin}(\lambda x)+\frac{\partial^2\textbf{}}{\partial x^2}e^{-\lambda x}\textbf{c}_2 \text{cos}(\lambda y)+\\
\space\\
+\frac{\partial^2\textbf{}}{\partial y^2}e^{-\lambda x}\textbf{c}_1\text{sin}(\lambda x)+\frac{\partial^2\textbf{}}{\partial y^2}e^{-\lambda x}\textbf{c}_2 \text{cos}(\lambda y)=\\
\space\\
=-2\textbf{c}_1\lambda^2e^{-\lambda x}\text{cos}(\lambda x)+\textbf{c}_2\lambda^2e^{-\lambda x}\text{cos}(\lambda y)+\\
\space\\
+0-\textbf{c}_2\lambda^2e^{-\lambda x}\text{cos}(\lambda y)=\\
\space\\
=-2\textbf{c}_1\lambda^2e^{-\lambda x}\text{cos}(\lambda x)."

We can't show what we were asked to show if we wrote the correct condition :(